# Replay Consolidation with Label Propagation for Continual Object Detection

### v2

@@ 3.2 Balanced Replay Memory
In the CLOD context, OCDM and Label Propagation work in synergy to optimize the replay memory, as demonstrated by the conducted ablation study of Sec. 5.3.
@@ 4.3 CL strategies details
Fine-Tuning 36.4 16.5 16.4 2.9 13.6 2.9
Replay [28] 54.5 60.8 56.9 40.0 14.1 8.0
@@ 5 Results
In this section, we discuss the outcomes of each method on the VOC and COCO CLOD benchmarks. Table 1 reports the results after all tasks have been completed for YOLOv8n. Additional results for YOLOv8M are provided in Sec. A.2. In Sec. 5.1, we discuss the results obtained in the 2-tasks scenarios, namely 15p5, 10p10, 19p1, and 40p40. In Sec. 5.2, we consider the more challenging scenarios 15p1 and 40p10.
@@ 5.2 Scenarios VOC 15p1 and COCO 40p10
As highlighted in Tab 1, Pseudo-Label works very well in a complex scenario like COCO 40p10. This is partially due to the very high intersection between tasks in terms of images, which allows the model to maintain knowledge of previous classes as it implicitly revisits them in the new task. However, in cases where the intersection between tasks is very low the Pseudo-Label performance is expected to be much worse. This is demonstrated by the VOC 15p1 scenario, where performance is low as only one new class is shown, and the intersection with previous tasks is much lower compared to COCO 40p10, where ten new classes are added each time.
Therefore, in practice, we have this interesting duality aspect where Replay for VOC 15p1 performs better than Pseudo-Label and vice versa for the COCO 40p10 scenario. Unlike Pseudo-label, interference between tasks in Replay is a problem since it implies saving the same image in memory several times with different labels, causing several issues, including an under-exploitation of memory. OCDM, despite having the same problem of missing annotations, is rewarded by carefully selecting the samples to save, which improves the final performance from 8 for Replay to 11 for OCDM in the COCO 40p10.
Then RCLPOD, which enhances the replay memory by improving the ground truth associated with the samples, receives a significant boost, achieving 20 in the COCO 40p10. This proves how much help the correct management of the information of the samples stored in the replay memory can be effective. Finally, even if LwF reaches good performance in both scenarios, the same problem highlighted for shorter tasks, namely, that the model achieves poor performance for new tasks is observed.
5.3 Ablation study
Replay OCDM LP FD Mask mAP
Model 1 14.1
Model 2 16.7
Model 3 19.7
Model 4 20.1
Model 5 24.1
Model 6 22.7
Table 2: Ablation study performed to study the effect of each component in RCLPOD, tested on COCO 40p10. LP is the Label Propagation mechanism, OCDM indicates the selection mechanism, FD represents the Feature Distillation component, and Mask represents the masking component.
@@ 6 Conclusions and Future Work
To improve the performances of Continual Object Detection, we propose a novel approach called RCPLOD. Our method, enhances the replay memory through class distribution balancing and Label Propagation, using existing data more effectively while solving the missing annotations and the task interference problem. RCLPOD outperforms existing techniques on well-established benchmarks such as VOC and COC.
@@
As future research, we believe that our approach can be used as a base for the proposal of new advancements in replay-based approaches. Indeed, the Label Propagation mechanism tested in the CLOD scenario proves its efficacy, making it a suitable base to build upon, since it facilitates more stable and memory-efficient models. In particular, the base of our strategy is designed to work independently of specific detector architectures. The technique manages memory and labeling in ways that are generally compatible with any object detection model using replay-based continual learning, simplifying the extension to other architectures. In addition, future research should give more attention to longer and more complex streams like COCO40p10 and beyond to better represent the real performances of the CL techniques in the CLOD setting. Following this direction, streams closer to realistic scenarios should be considered to represent better the challenges of Continual Object Detection.
@@ A.1 Further training hyper-parameters details
A.2 Results with YOLOv8m
As shown in Tab. 1, Pseudo-label performs slightly better than RCLPOD. However, as shown in 4, Pseudo-label performs worse than RCLPOD in terms of forgetting, while it outperforms the latter by plasticity.
As proved in [38], robustness to forgetting improves with scale of model size. Therefore, we also investigate the effect of the model size in the COCO 40p40 scenario. As presented in 4, RCLPOD performs better than Pseudo-Label, and the gap in terms of plasticity is lower than in the case of YOLOv8n. This fact shows that once the model size increases, the intrinsic plasticity of the model allows one to acquire new knowledge independently by the CL strategy applied to prevent forgetting.
A.3 Label Distribution for each VOC and COCO scenario
@@
A.4 LwF and stability-plasticity dilemma
@@ A.5 Memory efficiency
A.6 Comparison with literature
@@
A.7 Optimizing Class Distribution in Memory (OCDM)

### v3

@@ 3.2 Balanced Replay Memory
In the CLOD context, OCDM and Label Propagation work in synergy to optimize the replay memory, as demonstrated by the conducted ablation study of Sec. A.2.
@@ 4.3 CL strategies details
Fine-Tuning 36.4 16.5 16.4 3.0 13.6 2.9
Replay [28] 54.5 60.8 56.9 39.0 14.1 8.3
@@ 5 Results
In this section, we discuss the outcomes of each method on the VOC and COCO CLOD benchmarks. Table 1 reports the results after all tasks have been completed for YOLOv8n. Additional results for YOLOv8M are provided in Sec. A.3. In Sec. 5.1, we discuss the results obtained in the 2-tasks scenarios, namely 15p5, 10p10, 19p1, and 40p40. In Sec. 5.2, we consider the more challenging scenarios 15p1 and 40p10. In Sec. 5.3, we compare the different methods in terms of stability-plasticity.
@@ 5.2 Scenarios VOC 15p1 and COCO 40p10
As highlighted in Tab 1, Pseudo-Label works very well in a complex scenario like COCO 40p10. This is partially due to the very high intersection between tasks in terms of images, which allows the model to maintain knowledge of previous classes as it implicitly revisits them in the new task. However, in cases where the intersection between tasks is very low, the Pseudo-Label performance is expected to be much worse. This is demonstrated by the VOC 15p1 scenario, where performance is low as only one new class is shown, and the intersection with previous tasks is much lower compared to COCO 40p10, where ten new classes are added each time.
Therefore, in practice, we have this interesting duality aspect where Replay for VOC 15p1 performs better than Pseudo-Label and vice versa for the COCO 40p10 scenario. Unlike Pseudo-label, interference between tasks in Replay is a problem because it implies saving the same image in memory several times with different labels, which causes several issues, including an underexploitation of memory. OCDM, despite having the same problem of missing annotations, is rewarded by carefully selecting the samples to save, which improves the final performance from 8 for Replay to 11 for OCDM in the COCO 40p10.
Then RCLPOD, which enhances the replay memory by improving the ground truth associated with the samples, receives a significant boost, achieving 20 in the COCO 40p10. This proves how effective the correct management of the information about the samples stored in the replay memory can be. Finally, even if LwF reaches good performance in both scenarios, the same problem highlighted for shorter tasks, namely, that the model achieves poor performance for new tasks, is observed.
5.3 Stability-Plasticity Balance
In Tab. 2 we report a comparison in terms of stability and plasticity in the two more challenging scenarios. In particular, at each new task, we measure the mAP for old classes, namely classes that appear in the old task, and the mAP for new classes; at the end of the last task, we compute the average both for old classes and new classes, and we report the results in Tab. 2. By doing so, we measure the average performance of any method both in terms of forgetting (old) and plasticity (new). As highlighted by the results, our method is prone to reducing forgetting, and in the COCO 40p10 scenario, it is the best method for balancing both stability and plasticity. However, the VOC15p1 scenario shows there is still room for improvement in terms of plasticity. Regarding LwF, we discuss additional results regarding the stability-plasticity trade-off in A.5.
CL Method VOC 15p1 (mAP50) COCO 40p10 (mAP50-95)
old new all old new all
Fine-Tuning 9.0 38.3 3.0 0.0 28.7 2.9
Replay 47.0 53.8 39.0 12.7 11.9 8.3
LwF 48.8 31.6 41.5 20.5 15.4 16.2
OCDM 53.4 54.9 49.6 14.1 16.5 11.0
Pseudo-label 32.9 41.3 16.3 16.3 26.4 16.2
RCLPOD (Ours) 62.6 43.0 56.1 21.6 22.6 20.0
Table 2: Results for YOLOv8n for VOC 15p1 and COCO 40p10 scenarios. "old" is the average mAP for the old classes over all the tasks, "new" is the average mAP for the new classes over all the tasks.
@@ 6 Conclusions and Future Work
To improve the performance of Continual Object Detection, we propose a novel approach called RCPLOD. Our method enhances the replay memory through class distribution balancing and Label Propagation, using existing data more effectively while solving the missing annotations and the task interference problem. RCLPOD outperforms existing techniques on well-established benchmarks such as VOC and COC.
@@
As future research, we believe that our approach can be used as a base for the proposal of new advancements in replay-based approaches. Indeed, the Label Propagation mechanism tested in the CLOD scenario for the first time proves its efficacy, making it a suitable base to build upon since it facilitates more stable and memory-efficient models. In particular, the base of our strategy is designed to work independently of specific detector architectures. The technique manages memory and labeling in ways that are generally compatible with any object detection model using replay-based continual learning, simplifying the extension to other architectures. In addition, future research should give more attention to longer and more complex streams like COCO40p10 and beyond to better represent the real performances of the CL techniques in the CLOD setting. Following this direction streams closer to realistic scenarios should be considered to better represent the challenges of Continual Object Detection.
@@ A.1 Further training hyper-parameters details
A.2 Ablation study
Replay OCDM LP FD Mask mAP
Model 1 14.1
Model 2 16.7
Model 3 19.7
Model 4 20.1
Model 5 24.1
Model 6 22.7
Table 4: Ablation study performed to study the effect of each component in RCLPOD, tested on COCO 40p40. LP is the Label Propagation mechanism, OCDM indicates the selection mechanism, FD represents the Feature Distillation component, and Mask represents the masking component.
A.3 Results with YOLOv8m
As shown in Tab. 1, Pseudo-label performs slightly better than RCLPOD. However, as shown in 5, Pseudo-label performs worse than RCLPOD in terms of forgetting, while it outperforms the latter by plasticity.
As proved in [38], robustness to forgetting improves with scale of model size. Therefore, we also investigate the effect of the model size in the COCO 40p40 scenario. As presented in 5, RCLPOD performs better than Pseudo-Label, and the gap in terms of plasticity is lower than in the case of YOLOv8n. This fact shows that once the model size increases, the intrinsic plasticity of the model allows one to acquire new knowledge independently by the CL strategy applied to prevent forgetting.
A.4 Label Distribution for each VOC and COCO scenario
@@
A.5 LwF and stability-plasticity dilemma
@@ A.6 Memory efficiency
A.7 Comparison with literature
@@
A.8 Optimizing Class Distribution in Memory (OCDM)